Find the p-value for the hypothesis test. A random sample of size 51 is taken. The sample has a mean of 423 and a standard deviation of 84.

H0: µ = 400

Ha: µ ≠ 400

The p-value for the hypothesis test is .

Your answer should be rounded to 4 decimal places.
Answer #1

Solution :

\mu= 400

\bar x = 423

s = 84

n = 51

This is the two tailed test .

The null and alternative hypothesis is ,

H0 :  \mu  = 400

Ha : \mu    \neq 400

Test statistic = t

= (\bar x - \mu ) / s / \sqrt n

= (423 -400 ) /84 / \sqrt 51

= 1.955

Test statistic = t = 1.955

P-value = 0.0561